OpenStudy (anonymous):

WRITE AN EQUATION OF A LINE IN STANDARD FORM WITH THE GIVEN SLOPE THROUGH THE GIVEN POINT -2/3(6,9)

OpenStudy (unklerhaukus):

the standard form of a straight line is y= mx+b where m is the gradient, and b the y intercept (ie when x=0, y=b) m = -2/3, every point on the line satisfies this formula, so we can use a known point to solve for b the point (x,y) = (6,9) in this case so we can write x = 6, y=9 substituting the gradient and the point into the formula give us.. 9=(-2/3)6+b solving for b 9=-4+b b=13 With the gradient and the intercept y=(-2/3)x+13